OpenStudy (anonymous):

The surface area of a sphere is 324pi m². What is the volume of the sphere? Can someone please help me solve, I don't understand

OpenStudy (anonymous):

you must find the radius from the surface area of the sphere by setting it up like an equation like 324pi m= 4pi r^2 divide 324 by 4 you get 81 then find the square root of 81 which is 9 that would be your radius then plug in 9 as your radius for the volume of the sphere. V=(4/3)pi times 9 to the third power give you V=(4/3)pi x 729 divide by (4/3) give you 546.75pi would be your answer

OpenStudy (anonymous):

you must find the radius from the surface area of the sphere by setting it up like an equation like 324pi m= 4pi r^2 divide 324 by 4 you get 81 then find the square root of 81 which is 9 that would be your radius then plug in 9 as your radius for the volume of the sphere. V=(4/3)pi times 9 to the third power give you V=(4/3)pi x 729 divide by (4/3) give you 546.75pi would be your answer